# A Pairwise Comparison Relation-assisted Multi-objective Evolutionary Neural Architecture Search Method with Multi-population Mechanism

### v1

@@ 15      $t\leftarrow t+1$
III The Proposed Method For Evolutionary Multi-objective Neural Architecture Search
@@ III-A Overall Framework
An overview of the proposed algorithm is illustrated in Fig. 1. As shown in Fig. 1, our algorithm generally follows the basic process of genetic algorithm. Compared to existing ENAS methods, it mainly differs in the following two aspects: one is that a surrogate model is constructed to assist in the evaluation. Another is to add a multi-population mechanism into the process of generating new individuals.
Algorithm 1 shows the pseudo code of the proposed algorithm. Firstly, in order to construct an efficient surrogate model, some evaluated architectures are needed to construct the training dataset. So before the search process starts, an empty archive $\mathcal{A}$ to store training samples should be initialized (line 3). $N$ individuals are randomly sampled from the search space and then decoded to train. As mentioned above, we employ the weight sharing during the training. The weights inherited from the trained supernet are used as a warm up for the stochastic gradient descent (SGD) algorithm to improve the search efficiency (lines 4-9). Afterwards, we enter the main search loop of the algorithm (line 10). The surrogate model is the first to be constructed using the archive $\mathcal{A}$ (line 11). A detailed description of the surrogate model is given later in subsection III-C. Then we define a set to store individuals to be evaluated. Next, we use the proposed multi-objective evolutionary algorithm with multi-population mechanism (MP-MOEA for short) combine with our surrogate model to optimize both accuracy (Acc) and MAdds, i.e., maximize classification accuracy and minimize model complexity (line 12). This process is described in detail later in subsection III-D. To improve the sample efficiency of our search, we train the surrogate model using an online learning approach. The retained individuals of each generation are decoded and trained, and these retained offspring are used as new training samples to update the surrogate model (line 14-18). The above steps are repeated until the conditions for the end of the procedure are satisfied. Finally, after the multi-objective evolutionary search, the optimal architectures are selected from the pool of architectures based on the non-dominated sorting (line 21).
@@ III-B Search Space and Encoding
The search process of NAS starts with constructing a search space that can contain most CNNs for image tasks. In this paper, the backbone structure is based on MobileNetV3 [44]. The structure consists of three stages. The start stage extracts features, and the final stage outputs categories. These two parts do not need to be searched. The middle stage needs to be searched and consists of five sequentially connected MBConvBlocks [43] that progressively decrease the feature map size and increase the number of channels. Each block is composed of a series of layers, and every layer adopts the inverted bottleneck structure [45]: first, a $1\times 1$ convolution, which is used to convert the input channel to the expansion channel; second, a depth-wise convolution, which is the expansion channel and contains the parameter stride; at last, a $1\times 1$ convolution, which is used to convert from the expansion channel to the output channel. In addition, only the first layer uses stride 2 if the feature map size decreases and all the other layers in the block use stride 1. The number of layers of each convolution block (depth) is selected from $\left\{2,3,4\right\}$; for each layer, we search the expansion rate of the first $1\times 1$ convolution and the kernel size of the depth-wise separable convolution. The expansion rate is selected from $\left\{3,4,6\right\}$, and the kernel size is selected from $\left\{3,5,7\right\}$. Moreover, we also allow the CNNs to obtain multiple input image size (resolution), ranging from 192 to 256 with a stride of 4. For the encoding strategy, we use a fixed 46-bit integer string. If the encoding length of the architecture with fewer layers is less than 46 bits, the fixed length is achieved by padding with zeros. Fig. 2 shows the encoding strategy, and Fig. 3 shows the search space of this paper. As shown in these figures, the architecture is composed of five blocks. The encoding of this architecture is composed of image resolution and other parts representing five blocks. For each block, its encoding consists of the number of layers, the expansion rate, and the kernel size of each layer. In addition, the values in the encoding are all the index of considered options. As depicted in Fig. 2, the first bit “0” denotes the index of the first candidate resolution which means the resolution is 192. The remaining 45 bits represent five blocks, each of which is represented by 9 bits. e.g., the Block 1 can be encoded as “211200012”. The first bit “2” denotes the index of candidate layers which means the Block 1 has four layers. Then, the next four bits “1120” denotes the kernel size of each layer. Specifically, the kernel size in each layer in order is $\left\{5,5,7,3\right\}$. Similarly, the rest four bits denotes that the expansion rate in each layer in order is $\left\{3,3,4,6\right\}$. In Block 5, the gray bit “0” represents no layer and the encoding is padded with zero, which is not from considered options.
@@ III-C Surrogate Model
Unlike most of the existing surrogate-assisted methods used to predict the classification accuracy of architectures, we use the surrogate model to obtain the performance ranking of candidate architectures. In the evolutionary algorithm, the selection operation needs to be carried out according to the fitness values of individuals, and individuals with high fitness will be selected for the subsequent evolution. Therefore, it is more important to obtain the ranking of the candidates than to predict the true accuracy. We propose a surrogate model based on pairwise comparison relations, which is constructed on the basis of a classification model. Its inputs are vectors of the encoding combinations of pairwise architectures, and its outputs are labels indicating which one is the better solution. In this paper, we try to study the comparison relation between two individuals and transform the comparison relation learning problem into a binary classification problem. Binary classification models can be trained better with fewer samples. For example, $n(n-1)/2$ training samples can be obtained from a dataset containing $n$ labelled samples. Given a small number of samples, the surrogate model with better performance can be obtained through the extended training samples.
@@ III-D Multi-objective Evolutionary Algorithm Based on Multi-population Mechanism
In multi-objective evolutionary algorithms, diversity needs to be considered as well as convergence. During the practical evolution process, it is easy to fall into the local optimum without exploring the region of the optimal solution. To overcome this problem, we propose a multi-objective evolutionary algorithm based on the multi-population mechanism, see Algorithm 2 for details. The multi-population mechanism is mainly used in the selection operation to select parents for the next generation during the evolution process. Firstly, two sub-populations are constructed from the initial population $P_{t}$ (line 1). The main population is $E_{t}$, and the vice population is $F_{t}$. Specifically, the initial population $P_{t}$ is non-dominated sorted according to the predicted architectural rankings obtained from the above surrogate model and the calculated MAdds. After that, the non-dominated solutions of the first level are taken as the main population $E_{t}$ (line 2-5). Then, the individuals in the population $E_{t}$ are removed from the population $P_{t}$, and the crowding distances of the remaining solutions are calculated. $K$ individuals with high crowding ranking are selected to form the vice population $F_{t}$ (line 6). We set a threshold based on the number of generations, and when this threshold is less than $\theta$, Both parent $p_{1}$ and $p_{2}$ come from the main population $E_{t}$ (line 13). If this threshold is greater than $\theta$, parent $p_{1}$ comes from the primary population and parent $p_{2}$ comes from the secondary population (line 15). A number of offspring architectures are generated until the number of offspring is smaller than $m$ in the current generation. $m$ is a number used to limit the number of offspring. The threshold is a value related to the evolution generation, which is generated by the following equation:
@@
where $\delta$ is a randomly generated number between $[0,1]$, $g$ is the number of generation, and $G$ is the total number of generations.
@@ 2 NonDominatedSort($P_{t}$) according to the predicted ranking and complexity
10             if $\theta<Threshold$ then
@@ 27       $g\leftarrow g+1$
TABLE I: Comparison with state-of-the-art image classifiers on the CIFAR-10 dataset. The multi-objectives used for architecture optimization are performance and model complexity. *: MAdds computed using the genotypes provided by the authors. $\dagger$ denotes results from the reimplementation. SMEM-NAS(w/o): Architectures found by SMEM-NAS without the multi-population mechanism. We have bolded the minimum MAdds and maximum accuracy of the searched architectures.
Architecture Accuracy (%) MAdds (M) Search Cost (GPU Days) Search Method
EfficientNet-B0 [12] 98.1 387 - manual
MobileNetV2 [45] 95.74 300 - manual
PC-DARTS [46] 97.5 532* 0.3 GD
P-DARTS [47] 97.43 558* 0.1 GD
MixNet-M [20] $97.9^{\dagger}$ 359 - GD
NoisyDARTS [48] 97.63 534 0.4 GD
FairDARTS [49] 97.46 373 0.25 GD
sharpDARTS [50] 97.71 357 1.8 GD
AmoebaNet-B [51] 97.5 555 3150 EA
FairNAS-A [52] 98.2 391 12 EA
FairNAS-B [52] 98.1 348 12 EA
FairNAS-C [52] 98.0 324 12 EA
CARS [35] 97.43 728 0.4 EA
NSGA-Net [28] 97.98 817 4 EA
NASNet-A [6] 97.35 $608^{\ast}$ 1800 RL
SMEMNAS(w/o) 97.85 755 1.25 EA
SMEMNAS-S(ours) 97.96 200 1.25 EA
SMEMNAS-M(ours) 98.03 259 1.25 EA
SMEMNAS-L(ours) 98.13 305 1.25 EA
TABLE II: Comparison with state-of-the-art image classifiers on the CIFAR-100 dataset. $\dagger$ denotes results from the reimplementation. SMEM-NAS(transfer): the architecture, SMEMNAS-L, which searched on CIFAR-10.
Architecture Accuracy (%) MAdds (M) Search Cost (GPU Days) Search Method
EfficientNet-B0 [12] 88.1 400 - manual
MobileNetV2 [45] 80.8 300 - manual
PC-DARTS [46] 83.1 - 0.3 GD
MixNet-M [20] $87.1^{\dagger}$ 359 - GD
FairNAS-A [52] 87.3 391 12 EA
FairNAS-B [52] 87 348 12 EA
FairNAS-C [52] 86.7 324 12 EA
NSGA-Net [28] 85.58 817 4 EA
ZenNet [53] 84.4 487 0.5 EA
MUXNet-M [54] 86.11 400 11 EA
NASNet-A Large [6] 86.7 12031 1800 RL
NASNet-A mobile [6] 83.9 600 1800 RL
SMEMNAS(w/o) 86.65 348 1.25 EA
SMEMNAS(transfer) 87.18 305 1.25 EA
SMEMNAS-S(ours) 86.71 197 1.25 EA
SMEMNAS-M(ours) 87.25 250 1.25 EA
SMEMNAS-L(ours) 87.95 755 1.25 EA
@@ IV Experiments
In this section, we conduct a series of experiments to validate the effectiveness of our proposed method SMEM-NAS. Specifically, we firstly show our experimental configurations in Section IV-A. Next, Section IV-B presents and analyses our experimental results on benchmark datasets. Then we evaluate the effectiveness of the multi-population mechanism in Section IV-C. The comparison results with other algorithms in terms of accuracy and complexity, search cost is also discussed. Finally, Section IV-D show some ablation experiments for surrogate models and the initial population.
@@ IV-A Experimental Configurations
Regarding the benchmark datasets used in the experiments, we use CIFAR-10, CIFAR-100 and ImageNet datasets consistent with the existing ENAS works. In this study, the process of searching for architectures is performed separately on these three datasets. The performance of the final searched CNN architectures is also evaluated on three datasets. We evaluate the effectiveness of different architectures in terms of both classification accuracy and computational complexity.
@@
For each dataset, we start with 100 randomly sampled architectures from the search space as the initial population. Then the evolutionary search is performed, running for a total of 25 iterations. Except for the initial population, for each subsequent iteration, we retain and evaluate eight architectures from generated offspring. They are used for the training of the surrogate model and for the next iteration. So, in the whole search process, we evaluate 300 architectures in total, which are also used to train the surrogate model. Because OnceForAll is based on the ImageNet dataset, we fine-tune the weights inherited from the supernet for five epochs when searching on CIFAR-10 and CIFAR-100. All our experiments are performed on a single Nvidia RTX 3090 GPU card. Finally, we select three promising architectures from the obtained non-dominated solutions and then train them for 200 epochs by the standard SGD optimizer with momentum, where the initial learning rate, the momentum rate and the batch size are set to 0.01, 0.9 and 128.

### v2

@@ 16      $t\leftarrow t+1$
III The Proposed Method of Evolutionary Multi-objective Neural Architecture Search
@@ III-A Overall Framework
An overview of the proposed algorithm is illustrated in Fig. 1. As shown in Fig. 1, our algorithm generally follows the basic process of genetic algorithm. Compared to existing NAS methods, the proposed method has two key innovations: one is that a surrogate model is constructed to assist the evaluation. Another is that a multi-population mechanism is added into the process of generating new individuals.
Algorithm 1 shows the pseudo code of the proposed algorithm. Firstly, in order to construct an efficient surrogate model, some evaluated architectures are needed to construct the training dataset. So before the search process starts, an empty archive $\mathcal{A}$ for storing training samples is initialized (line 3). $N$ individuals are randomly sampled from the search space and then are decoded for training. As mentioned above, we employ the weight sharing during the training. The weights inherited from the trained supernet are used as a warm up for the stochastic gradient descent (SGD) algorithm to improve the search efficiency (lines 4-10). Afterwards, there is a main search loop of in the algorithm (line 11). The surrogate model is constructed using the archive $\mathcal{A}$ (line 12). A detailed description of the surrogate model is given later in subsection III-C. Then we define a set to store individuals to be evaluated. Next, we use the proposed multi-objective evolutionary algorithm with multi-population mechanism (MP-MOEA for short) combine with our surrogate model to optimize both accuracy (Acc) and MAdds, i.e., maximize classification accuracy and minimize model complexity (line 13). This process is described in detail later in subsection III-D. To improve the sample efficiency of our search, we train the surrogate model using an online learning approach. The retained individuals of each generation are decoded and trained, and these retained offspring are used as new training samples to update the surrogate model (lines 15-19). The above steps are repeated until the conditions for the end of the procedure are satisfied. Finally, after the multi-objective evolutionary search, the optimal architectures are selected from the pool of architectures based on the non-dominated sorting (line 22).
@@ III-B Search Space and Encoding
The search process of NAS starts with constructing a search space that can contain most CNNs for image tasks. In this paper, the backbone structure is based on MobileNetV3 [52]. The structure consists of three stages. The start stage extracts features, and the final stage outputs categories. These two parts do not need to be searched. The middle stage needs to be searched and consists of five sequentially connected MBConvBlocks [51] that progressively decrease the feature map size and increase the number of channels. Each block is composed of a series of layers, and every layer adopts the inverted bottleneck structure [53]: first, a $1\times 1$ convolution, which is used to convert the input channel to the expansion channel; second, a depth-wise convolution, which is the expansion channel and contains the parameter stride; at last, a $1\times 1$ convolution, which is used to convert from the expansion channel to the output channel. The number of layers of each convolution block (depth) is selected from $\left\{2,3,4\right\}$; for each layer, we search the expansion rate of the first $1\times 1$ convolution and the kernel size of the depth-wise separable convolution. The expansion rate is selected from $\left\{3,4,6\right\}$, and the kernel size is selected from $\left\{3,5,7\right\}$. Moreover, we also allow the CNNs to obtain multiple input image size (resolution), ranging from 192 to 256 with a stride of 4. Therefore, there are approximately $10^{20}$ different architectures in the search space. For the encoding strategy, we use a fixed 46-bit integer string. If the encoding length of the architecture with fewer layers is less than 46 bits, the fixed length is achieved by padding with zeros. Fig. 2 shows the encoding strategy, and Fig. 3 shows the search space of this paper. As shown in these figures, the architecture is composed of five blocks. The encoding of this architecture is composed of image resolution and other parts representing five blocks. For each block, its encoding consists of the number of layers, the expansion rate, and the kernel size of each layer. In addition, the values in the encoding are all the index of considered options. As depicted in Fig. 2, the first bit “0” denotes the index of the first candidate resolution which means the resolution is 192. The remaining 45 bits represent five blocks, each of which is represented by 9 bits. e.g., the Block 1 can be encoded as “211200012”. The first bit “2” denotes the index of candidate layers which means the Block 1 has four layers. Then, the next four bits “1120” denote the kernel size of each layer. Specifically, the kernel size in each layer in order is $\left\{5,5,7,3\right\}$. Similarly, the rest four bits denote that the expansion rate in each layer in order is $\left\{3,3,4,6\right\}$. In Block 5, the gray bit “0” represents no layer and the encoding is padded with zeros, which is not from considered options.
@@ III-C Surrogate Model
Unlike most of the existing surrogate-assisted methods used to predict the classification accuracy of architectures, we use the surrogate model to obtain the performance ranking of candidate architectures. In the evolutionary algorithm, the selection operation needs to be carried out according to the fitness values of individuals, and individuals with high fitness will be selected for the subsequent evolution. Therefore, it is more important to obtain the ranking of the candidates than to predict the true accuracy. We propose a surrogate model based on pairwise comparison relations, which is constructed on the basis of a classification model. Its inputs are vectors of the encoding combinations of pairwise architectures, and its outputs are labels indicating which one is the better solution. In this paper, we try to study the comparison relation between two individuals and transform the comparison relation learning problem into a binary classification problem. Binary classification models can be trained better with fewer samples. For example, given $n$ labeled architectures, we generate training samples by comparing all unique pairs of architectures, and $\frac{n(n-1)}{2}$ training samples can be obtained. Given a small number of samples, the surrogate model with better performance can be obtained through the extended training samples.
@@ III-D Multi-objective Evolutionary Algorithm Based on Multi-population Mechanism
In multi-objective evolutionary algorithms, diversity needs to be considered as well as convergence. During the practical evolution process, it is easy to fall into the local optimum without exploring the region of the optimal solution. To overcome this problem, we propose a multi-objective evolutionary algorithm based on the multi-population mechanism, see Algorithm 2 for details. The multi-population mechanism is mainly used in the selection operation to select parents for the next generation during the evolution process. Firstly, two sub-populations are constructed from the initial population $P_{t}$ (line 1). The main population is $E_{t}$, and the vice population is $F_{t}$. Specifically, the initial population $P_{t}$ is non-dominated sorted according to the predicted architectural rankings obtained from the above surrogate model and the calculated MAdds. After that, the non-dominated solutions of the first level are taken as the main population $E_{t}$ (lines 2-5). Then, the individuals in the population $E_{t}$ are removed from the population $P_{t}$, and the crowding distances of the remaining solutions are calculated. $K$ individuals with high crowding ranking are selected to form the vice population $F_{t}$ (line 6). We set a threshold based on the number of generations, and when this threshold is less than $\theta$, Both parent $p_{1}$ and $p_{2}$ come from the main population $E_{t}$ (line 13). If this threshold is greater than $\theta$, parent $p_{1}$ comes from the primary population and parent $p_{2}$ comes from the secondary population (line 15). A number of offspring architectures are generated until the number of offspring is greater than or equal to $m$ in the current generation. $m$ is a number used to limit the number of offspring. The threshold is a value related to the evolution generation, which is generated by the following equation:
@@
$\delta=$$\left\{\begin{array}[]{lc}random(0,0.7)&g<\frac{1}{4}G\\ random(0,1)&g\geq\frac{1}{4}G\\ \end{array}\right.$ (7)
where $Threshold$ and $\delta$ are randomly generated in every generation, $g$ denotes the serial number of the current generation, and $G$ is the total number of generations.
@@ 2 NonDominatedSort($P_{t}$) according to the predicted ranking and complexity
10             if $Threshold\textgreater\theta$ then
@@ 27       $g\leftarrow g+1$
TABLE I: Comparison with state-of-the-art image classifiers on the CIFAR-10 dataset. The multi-objectives used for architecture optimization are performance and model complexity. *: MAdds computed using the genotypes provided by the authors. $\dagger$ denotes results from the reimplementation. SMEM-NAS(w/o): Architectures found by SMEM-NAS without the multi-population mechanism. Search cost: GPU days of the searching process.
Architecture
Top-1
MAdds
Params
Search
Cost
Device Year
Search
Method
EfficientNet-B0 [13] 98.1 387 4.0 - - 2019 manual
iDARTS [54] 97.75 - 3.6 0.2 Tesla V100 2023 GD
CDARTS [55] 97.52 - 3.9 0.3 Tesla V100 2023 GD
SWD-NAS [56] 97.49 519 3.17 0.13 RTX 3090 2024 GD
PA-DARTS [57] 97.59 578 3.75 0.36 RTX 3090 2024 GD
NoisyDARTS [58] 97.63 534 3.3 0.4 Tesla V100 2021 GD
GENAS [59] 97.55 504 3.53 0.26 RTX 3090 2024 GD
RelativeNAS [60] 97.66 - 3.93 0.4 GTX 1080Ti 2023 GD
DARTS-PT-CORE [61] 97.5 - 3.6 0.8 Tesla V100 2024 GD
NPENAS-NP [62] 97.46 - 3.5 1.8 RTX 2080Ti 2023 EA
EEEA-Net-C [63] 97.54 - 3.6 0.52 RTX 2080Ti 2021 EA
CGP-NAS [64] 96.3 636 4.04 11.5 RTX 1080Ti 2023 EA
SMCSO [65] 97.12 - 3.46 1.32 RTX 2080Ti 2025 EA
SPNAS [66] 98.2 - 6.33 1.4 RTX 3090 2025 EA
FairNAS-A [49] 98.2 391 - 12 Tesla V100 2021 EA
CARS [37] 97.38 728 3.6 0.4 Tesla V100 2020 EA
ESENet [67] 96.44 540 4.53 9 RTX 2080Ti 2023 EA
NSGANetV2 [25] 98.4 468 - - - 2020 EA
M2M-Net [68] 97.56 - 3.79 6 RTX 3080 2024 EA
NASNet-A [7] 97.35 $608^{\ast}$ 3.3 1800 Tesla P100 2018 RL
BNAS [69] 97.03 - 4.7 0.19 - 2022 RL
DBNAS [70] 97.58 - 2.4 0.9 RTX 3090 2025 RL
SMEMNAS-S(w/o) 97.51 186 4.7 1.25 RTX 3090 - EA
SMEMNAS-M(w/o) 97.78 313 5.8 1.25 RTX 3090 - EA
SMEMNAS-L(w/o) 97.85 755 6.4 1.25 RTX 3090 - EA
SMEMNAS-S(ours) 97.96 200 4.7 1.25 RTX 3090 - EA
SMEMNAS-M(ours) 98.03 259 5.7 1.25 RTX 3090 - EA
SMEMNAS-L(ours) 98.13 305 5.0 1.25 RTX 3090 - EA
TABLE II: Comparison with state-of-the-art image classifiers on the CIFAR-100 dataset. $\dagger$ denotes results from the reimplementation. SMEMNAS-transfer: the architecture, SMEMNAS-L, which searched on CIFAR-10.
Architecture
Top-1
MAdds
Params
Search
Cost
Device Year
Search
Method
EfficientNet-B0 [13] 88.1 400 4.0 - - 2019 manual
MobileNetV2 [53] 80.8 300 3.4 - - 2018 manual
PC-DARTS [71] 83.1 - 3.6 0.3 GTX 1080Ti 2019 GD
SWD-NAS [56] 83.87 598 3.56 0.13 RTX 3090 2024 GD
DARTS-PT-CORE [61] 83.27 - 3.6 0.8 Tesla V100 2024 GD
PA-DARTS [57] 83.62 578 3.75 0.36 RTX 3090 2024 GD
RelativeNAS [60] 84.14 - 3.98 0.4 GTX 1080Ti 2023 GD
GENAS [59] 83.14 504 3.53 0.26 RTX 3090 2024 GD
SMCSO [65] 80.66 - 3.72 2 RTX 2080Ti 2025 EA
FairNAS-A [49] 87.3 391 - 12 Tesla V100 2021 EA
CGP-NAS [64] 79.37 827 5.9 11.28 RTX 1080Ti 2023 EA
ESENet [67] 76.35 540 4.53 9 RTX 2080Ti 2023 EA
M2M-Net [68] 84.77 - 3.83 6 RTX 3080 2024 EA
SPNAS [66] 87.26 - 6.7 1.6 RTX 3090 2025 EA
NASNet-A Large [7] 86.7 12031 - 1800 Tesla P100 2018 RL
DBNAS [70] 85.18 - 2.5 0.9 RTX 3090 2025 RL
SMEMNAS-S(w/o) 86.05 194 3.5 1.25 RTX 3090 - EA
SMEMNAS-M(w/o) 86.4 242 4.1 1.25 RTX 3090 - EA
SMEMNAS-L(w/o) 86.65 348 5.8 1.25 RTX 3090 - EA
SMEMNAS-transfer 87.18 305 5.0 1.25 RTX 3090 - EA
SMEMNAS-S(ours) 86.71 197 3.9 1.25 RTX 3090 - EA
SMEMNAS-M(ours) 87.25 250 4.1 1.25 RTX 3090 - EA
SMEMNAS-L(ours) 87.95 755 6.4 1.25 RTX 3090 - EA
@@ IV Experiments
In this section, we conducted a series of experiments to validate the effectiveness of our proposed method SMEM-NAS. Specifically, we firstly show our experimental configurations in Section IV-A. Next, Section IV-B presents and analyses our experimental results on benchmark datasets. Then we evaluate the effectiveness of the multi-population mechanism in Section IV-C. The comparison results with other algorithms in terms of different perspectives are discussed, including accuracy, complexity, search cost, device and year of publication. Finally, Section IV-D shows some ablation experiments for surrogate models and the initial population.
@@ IV-A Experimental Configurations
Regarding the benchmark datasets used in the experiments, we used CIFAR-10, CIFAR-100 and ImageNet datasets consistent with the existing ENAS works. In this study, the process of searching for architectures was performed separately on these three datasets. The performance of the final searched CNN architectures was also evaluated on three datasets. We evaluated the effectiveness of different architectures in terms of both classification accuracy and computational complexity.
@@
For each dataset, we started with 100 randomly sampled architectures from the search space as the initial population. The evolutionary search process was executed for a total of 25 iterations. Except for the initial population, for each subsequent iteration, we retained and evaluated eight architectures from generated offspring. They were used for the training of the surrogate model and for the next iteration. So, in the whole search process, we evaluated 300 architectures in total, which were also used to train the surrogate model. Since evaluating architecture performance is time-time-consuming, we adopted the weight-sharing technique to accelerate the training process to improve the search efficiency. The OnceForAll [51] supernet is trained with the progressive shrinking algorithm, in which the largest network is fine-tuned to support sub-networks and they are added to the sampling space by sharing weights. During the search process, the architecture weights were inherited from the trained OnceForAll as initial weights for training. Because OnceForAll is based on the ImageNet dataset, we fine-tuned the weights inherited from the supernet for five epochs when searching on CIFAR-10 and CIFAR-100. All our experiments were performed on a single Nvidia RTX 3090 GPU card. Finally, we selected three promising architectures from the obtained non-dominated solutions and then trained them for 200 epochs by the standard SGD optimizer with momentum, where the initial learning rate, the momentum rate and the batch size were set to 0.01, 0.9 and 128. More detailed information can be found in our publicly available code https://github.com/ccz-enas/SMEM-NAS.
IV-B Results on Standard Datasets
In this section, we show the results of our method on benchmark datasets to validate its effectiveness. We conducted the search and trained the final architectures separately on three datasets, which is different from many existing NAS methods. They mostly adopt a transfer learning setup where architectures searched on one dataset are directly applied to others. We think the transferred architectures may not be optimal on another dataset. Then, in order to validate the effectiveness of the proposed method, we compare the non-dominated architectures obtained by SMEM-NAS with those obtained by other state-of-the-art NAS methods. The selected peer methods can be broadly divided into three categories: manually designed by human experts, EA-based, and non-EA-based (e.g., RL and GD). In addition, we give the searched results by SMEM-NAS without the multi-population mechanism.
Figure 5: Accuracy and number of multi-adds in millions on CIFAR-10. Models from multi-objective approaches are joined with lines.
Results on CIFAR-10: After completing the search using our method, we selected three promising architectures, and for ease of representation we named them SMEMNAS-S/M/L (sorted by MAdds). Fig. 5 and Table I show the results of SMEM-NAS and other methods on CIFAR-10. Fig. 5 visualizes the dominant relationship between the network models in terms of accuracy and model complexity. It is obvious that our models always outperform most other models. Table I shows more details. Firstly, from the perspective of time, our method takes 1.25 GPU days and significantly reduces the search time compared with methods based on EA and RL. The surrogate model we used effectively accelerates the search process. Compared with the GD-based method, although there are some deficiencies in time, the architecture we searched has better performance. Our model achieves the accuracy of 98.13%, surpassing almost all peer competitors. Among them, our model is slightly lower than FairNAS-A, but has nearly 100M lower MAdds and the search time is much less than FairNAS. In general, SMEM-NAS is superior to or consistent with other models in different metrics. Moreover, we also present the results obtained without employing the multipopulation mechanism (denoted as w/o), thereby demonstrating the effectiveness of this mechanism.
Results on CIFAR-100: Similar to the settings on CIFAR-10, we also conducted experiments on CIFAR-100. Fig. 6 and Table II show the results of our model and other methods. We also provide three models on this dataset, i.e., large, medium and small, in which the highest accuracy is 87.95%. From Fig. 6, it is clearly observed that our architectures dominate most methods in terms of accuracy and complexity. Table II shows more details. On this dataset, although our results are slightly inferior to EfficientNet, the hand-crafted model requires significant expertise and labour. Additionally, we can see that the architecture with the highest accuracy searched on the CIFAR-10 dataset transferred to CIFAR-100 does not perform optimally, achieving 87.18% accuracy. Moreover, the search time of our method on this dataset is shorter than that of all other methods that also employ evolutionary algorithms.